SPIN-ORBIT MIXING IN THE A^{1}Σ_{u}^{+} AND b^{3}Π_{0u} STATES OF THE RUBIDIUM DIMER

AMANDA ROSS, PATRICK CROZET, Université Lyon 1 and CNRS (LASIM, UMR 5579), 69622 Villeurbanne, France; \underline{HOUSSAM SALAMI}, THOMAS BERGEMAN, Department of Physics and Astronomy, SUNY, Stony Brook, NY 11794-3800; OLIVIER DULIEU, Laboratoire Aimé Cotton, CNRS, Bât. 505, Campus d’Orsay, 91405 Orsay, France; CHRISTIAN LISDAT, Physikalisch-Technische Bundesanstalt, Braunschweig, Germany; BEDIHA BESER, JIANMEI BAI, A. MAR-JATTA LYYRA and SVETLANA KOTOCHIGOVA, Department of Physics, Temple University, Philadelphia, PA 19122-6082.

The A^{1}Σ_{u}^{+} and b^{3}Π_{0u} states of the rubidium dimer has been re-investigated experimentally, adding to the considerable array of upper state term values determined from earlier work (performed at Laboratoire Aimé Cotton to characterize the electronic ground state). A polarization labelling experiment probed a series of levels from v=0, J=71 in the X^{1}Σ_{g}^{+} state, and (A^{1}Σ_{u}^{+} ∼ b^{3}Π_{0u}) → X^{1}Σ_{g}^{+} fluorescence spectra have been recorded on an FT spectrometer following excitation of low-lying vibrational levels in the A state by a Ti:sapphire laser operating with long wavelength optics. Data for ^{85}Rb_{2}, ^{85}Rb^{87}Rb and ^{87}Rb_{2} are modeled using the discrete variable representation, fitting to numerical potential curves and Morse-type spin-orbit functions starting from ab initio potentials and spin-orbit functions. The fit has confirmed absolute vibrational numbering in the A state (the vibrational assignment in the b state looks convincing but is not definitive). The fit currently returns a root mean square residual of 0.075 cm^{−1}, which is 15 × the estimated experimental uncertainty, reflecting the fact that information particularly on the b state is still sparse. Including the b^{3}Π_{1} component in the Hamiltonian did not improve the fit. Nevertheless, the fit provides useful estimates of term values and of spin-orbit mixing effects.

Work at Stony Brook was supported by NSF grant PHY 0652459 and at Temple University by NSF PHY 0555608.